OpenStudy (anonymous):

what is 2.66 in fraction form?

OpenStudy (diyadiya):

266/100

OpenStudy (anonymous):

133/50

OpenStudy (diyadiya):

Yes simplified^^^

OpenStudy (anonymous):

No like 2 and what over what

OpenStudy (anonymous):

2 and 33/50

OpenStudy (anonymous):

what is 33/50 reduced down to

OpenStudy (anonymous):

It can't be reduced. If you want the decimal form it is .66. Because 2+.66=2.66 as you asked.

OpenStudy (anonymous):

well we have a puzzle and that 2 33/50 is not one of the answers

OpenStudy (diyadiya):

what are the choices?

OpenStudy (anonymous):

there are a lot of anwers

OpenStudy (anonymous):

anwser*

OpenStudy (anonymous):

the problem is 16= -8 - 9y find y

OpenStudy (diyadiya):

add 8 on both sides

OpenStudy (diyadiya):

and then divide by -9

OpenStudy (anonymous):

ok did that

OpenStudy (anonymous):

see thats where I think I'm getting the wrong answer

OpenStudy (diyadiya):

16=-8-9y 16+8=-8+8-9y 24=-9y -9y=24 y=-24/9= - 8/3

OpenStudy (diyadiya):

right?

OpenStudy (diyadiya):

\[\frac{-8}{3}= -2 \frac{2}{3}\]

OpenStudy (anonymous):

Oh ok. Thank you. I'm not good w/ fractions

OpenStudy (diyadiya):